#1 mks R aML VIN M + _M- Vo FlalD 25KA #2 20K 20KA ww 4V FIND VO
Answer #1

R3 R5 VIN IN Vp Muz Va Es Uni ve MM- Find the vo VIN Note; Assume op-amp is ideal ( ENSURE OA and vp = v) → Taking kc) at ModT va = Ro Vw + Rz No - Va = Rom vin + Revo. O - Taking kel at Mode Ver let is = En, to (izon) MMI + Na-VN = R6 R5 UN, - No ReVo - Rz Ky vo R2 R4 Vint RIRS R3 Rs R2 R3 R4 RC Rs R5 - Rz R4) # 2 25kw took MM- ROK 66V Qok 20kw Slokn 2V ) 5 Find the Vai-=latin (FOR |Vpzov ve-VM 46- VA 2010 V - Na = VNTOV Hok 4 Va = 40 Va= - 8V Redraw the above cinmit 25kw to look M . 20kr Los6- VW -8- VN Q- VN VW-Vo. 25k 20k lok look | Us=ov) - Vo + S + Źo = 816) – 8(10) + 20 (2) = 200 48-50 +40 to -VO lo- - Vo = -